Hero Manufacturing has 8.9 million shares of common stock outstanding. The current share price is $73 and the book value per share is $6. The company also has two bond issues outstanding. The first bond issue has a face value of $65 million, a coupon rate of 6.3 percent and sells for 109.3 percent of par. The second issue has a face value of $52.8 million, a coupon rate of 8.2 percent and sells for 113.1 percent of par. The first issue matures in 10 years, the second in 26 years.

Suppose the company’s stock has a beta of 1.4. The risk-free rate is 3.2 percent and the market risk premium is 7.3 percent. Assume that the overall cost of debt is the weighted average implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 21 percent.

What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

WACC = We * Ke + Wd1 * Kd1 + Wd2 * Kd2 where

We , Wd1, Wd2 = Market value weights of Equity, debt 1 and debt 2 respectively

Ke, Kd1, Kd2 = Cost of Equity, debt 1 and debt 2 respectively

Market value weights

Product Market value Weights
Equity (8900000*73) 649700000 649700000/780461800 = 0.832455861388732
First Bond (65000000*109.3%) 71045000 71045000/780461800 = 0.0910294392371286
Second Bond (52800000*113.1%) 59716800 59716800/780461800 = 0.0765146993741398
Total 780461800 1.00

Cost of Equity

Ke = Risk free rate + beta * market risk premium

= 3.2% + 1.4 * 7.3% = 13.42%

Cost of First bond

Kd1 = YTMannual of bond 1 where YTM is yield to maturity

YTMsemiannual = YTM approximate C+(FP - MP)/n (FP+ MP )/2

C = semiannual coupon = 65000000 * 6.3% * 6/12 = 2047500

FP = face value = 65000000

MP = market price = 65000000*109.3% = 71045000

n = number of payments = 10 * 2 = 20

YTMsemiannual = (2047500 + ((65000000 - 71045000)/20)) / ((65000000 + 71045000)/2)

= (2047500 + ((-302250) / 68022500

= 1745250 / 68022500 = 0.0256569517439083

YTMannual = (1+YTMsemiannual)^2 - 1 = ((1+0.0256569517439083)^2)-1 = 5.19721826606057%

Kd1 = 5.19721826606057%

Cost of Second bond

C = semiannual coupon = 52800000 * 8.2% * 6/12 = 2164800

FP = face value = 52800000

MP = market price = 52800000*113.1% = 59716800

n = number of payments = 26* 2 = 52

YTMsemiannual = (2164800 + ((52800000 - 59716800)/52)) / ((52800000 + 59716800)/2)

= (2164800 + (-133015) / 56258400

= 2031785 / 56258400 = 0.0361152221781035

Kd2 = ((1+0.0361152221781035)^2 - 1) = 7.3534753629181%

Now that we have all the required variables we can calculate WACC

WACC = (0.832455861388732 * 13.42%) + (0.0910294392371286 * 5.19721826606057%) + (0.0765146993741398 * 7.3534753629181%)

= 11.1715576598368% + 0.473099864352455% + 0.562648956748822% = 12.21%

WACC = 12.21